Write the standard form of the equation of the line that passes through the given points
(-1, -2), (0, 3)

Answer by checkley71(8403)

FIRST YOU NEED TO FIND THE SLOPE (m) (Y2-Y1)/(X2-X1)
(3+2)/(0+1)
5/1=5 FOR THE SLOPE
NOW REPLACE ONE SET THE X & Y VALUES IN THE LINE EQUATION & SOLVE FOR (b) THE Y INTERCEPT.
Y=mX+b
-2=5*-1+b
-2=-5+b
b=-2+5
b=3 FOR THE Y INTERCEPT THUS THE LINE EQUATION IS:
Y=5X+3